### v1

@@ 4 Software: The challenge of increasing fragmentation
The exascale era has reiterated the need for a more unified and standardized software stack for GPU programming. The fragmentation creates barriers to code portability, hinders performance optimization, and increases the development effort required for scientific applications. As we move forward, the community must address these challenges. Initiatives like SYCL, OpenMP, and community-driven efforts like Kokkos and RAJA look promising toward achieving a more unified programming experience. However, vendors, developers, and researchers must collaborate closely to establish common standards and standardize a more cohesive and interoperable software ecosystem for GPU computing.
5 QM computation in the current HPC landscape
Quantum mechanical (QM) methods encompass a broad spectrum of applications, each with varying computational demands and levels of accuracy. As depicted in Fig. 3, density functional theory (DFT) represents a computationally efficient approach, offering a favorable balance between precision and performance, making it a popular choice for many systems. On the other end of the spectrum, wave-function-based methods are indispensable when higher accuracy is required to capture the intricate properties of more complex systems. Quantum Monte Carlo (QMC) methods also offer a path to high precision for larger systems by employing stochastic techniques to solve the Schrödinger equation.
The nominal computational cost of these methods varies significantly. DFT is generally efficient, scaling with the number of electrons as , primarily due to the diagonalization of the Kohn-Sham eigenvalue problem. In contrast, wave-function-based methods such as configuration interaction (CI), density matrix renormalization group (DMRG), and coupled-cluster (CC) methods demand considerably more computational resources, with scaling complexities that increase steeply depending on the method’s sophistication. For example, the computational cost of the “gold standard” in quantum chemistry, coupled-cluster singles and doubles with perturbative triples (CCSD(T)), scales as . QMC methods, particularly in their commonly used real-space variants, occupy an intermediate position in terms of computational cost, with a scaling of per Monte Carlo step, largely governed by the repeated evaluation of the Slater determinant.
@@ 5 QM computation in the current HPC landscape
Each QM method relies on distinct computational techniques (e.g., fast Fourier transforms, matrix-vector operations, diagonalization of sparse or dense matrices), which determine the scaling behavior, while the specific hardware implementation governs the prefactor associated with this scaling. For instance, matrix multiplication kernels are central to CC methods that utilize large tensors, enabling efficient use of tensor cores and resulting in a relatively small prefactor despite the high scaling complexity. On the other hand, although QMC methods are inherently massively parallel, they often operate on relatively small matrices, leading to a large prefactor; therefore, scaling these approaches on accelerators requires significant effort [55].
The computational physics and chemistry communities have implemented these and other QM approaches in various software packages, facilitating discoveries across multiple disciplines. Among the widely used electronic structure codes, ABINIT [56], BerkeleyGW [57], BigDFT [58], CP2K [59], FHI-aims [60], NWChemX [61], Q-Chem [62], Quantum ESPRESSO (QE) [63], and VASP [64] have been particularly successful in harnessing HPC advancements for QM calculations, each with unique strengths and target applications. These codes have made significant strides in simulating larger systems, incorporating advanced algorithmic and methodological developments, and enhancing parallel efficiency, particularly on pre-exascale architectures [29, 65].
@@
Recent efforts to adapt to the heterogeneous architectures of HPC systems highlight the critical importance of balancing performance with portability. Some codes, such as ABINIT and QE, have adopted hardware-agnostic offload programming models (OpenMP or OpenACC), while others, like CP2K, have opted for non-portable, kernel-based models (CUDA and HIP). BerkeleyGW provides both CUDA and OpenACC alternatives, with the latter achieving performance within ten percent of the former [57]. Additionally, different communities are actively developing specialized performance libraries to execute computationally intensive tasks on accelerators. Examples include the Distributed Block Compressed Sparse Row Matrix (DBCSR) library [66] for NVIDIA and AMD GPUs and the LibintX [67, 68] library, which provides efficient implementations of Gaussian integral evaluation, central to integral-direct implementations. The QMCkl library, developed within the TREX CoE, offers high-performance computation kernels for QMC simulations using OpenMP, OpenACC, and SYCL [69]. The wave-function community [70, 27, 61, 71, 72, 73] has been extensively developing tensor libraries that automatically parallelize and accelerate operations on multidimensional arrays, taking advantage of modern HPC architectures [74, 75, 76].
6 AI and quantum simulations
While both AI and QM simulations extensively need the power of modern computing, the nature of computations involved in each field and their respective challenges in reaching exascale performance are fundamentally different. AI, particularly machine learning (ML), has exhibited significant progress in recent years due to its ability to extract patterns and learn complex relationships from data [79, 80]. The success (see Fig. 4(4(a))) is mainly due to the availability of large datasets, advancements in algorithm development, and the accessibility of HPC resources, especially those equipped with GPUs and tensor processing units (TPUs).
The computations involved in AI training primarily revolve around optimizing the weights and biases of a neural network to minimize a specific loss function, which measures the difference between predicted and actual target values. Such tasks often involve reduced-precision matrix multiplications, convolutions, and activation function evaluations. Optimization techniques like stochastic gradient descent and its variants are also widely employed, involving iterative updates to network parameters based on gradients of the loss function calculated using backpropagation [81, 82]. These computations are inherently parallel, as the gradients for different data points can be computed independently and then averaged. Furthermore, model parallelism, where different parts of a model are handled by different processors, allows for efficient training of complex models with billions of parameters [83]. Figure 4(4(a)) illustrates the exponential growth in computational requirements for training large AI models, that double every 3.4 months compared to conventional Moore’s law [84]. The inherent parallelism in AI training makes it particularly well-suited for GPUs, which excel at executing the same operation on multiple data points simultaneously.
Consequently, the impressive acceleration of almost , illustrated in Table 1, for matrix multiplications using reduced precision on GPUs can be readily exploited by AI, while QM simulations typically require double precision where only a tenfold speedup can be achieved for matrix multiplications using vendor libraries. For kernels written in OpenMP, OpenACC, or OpenCL, lacking the tensor core boost, an acceleration factor of less than five is usually expected in QM simulations. In the current QM landscape, many algorithms exhibit low arithmetic intensity, largely due to their historical development and the limitations of earlier hardware. As a result, the potential advantages of GPUs are not always clear for these traditional algorithms. To fully harness the computational power of modern GPUs, particularly their ability to perform dense matrix multiplications at reduced precision, new algorithms must be developed. These new approaches would need to align more closely with GPU strengths, optimizing for parallelism and reduced precision where possible, rather than relying on the double-precision calculations and sparse matrix operations that have characterized QM methods to date.
Interestingly, scientists have started to utilize the power of AI/ML to address the accuracy-efficiency dilemma of quantum mechanical calculations. For example, machine learning potentials, which are generally trained on ab initio data, can predict potential energy surfaces and forces with the accuracy of the underlying quantum method, but at a much lower computational cost [85]. An efficient materials discovery and design using high-throughput calculations is now possible [86] with public datasets [87] and efficient optimization techniques [88]. Furthermore, neural network wavefunctions in combination with quantum Monte Carlo methods have shown promise in achieving highly accurate solutions of the electronic Schrödinger equation for electronic systems [89, 81, 82, 90, 91].
7 Outlook and conclusions
@@
The future of quantum mechanical simulations at the exascale level hinges on effectively navigating the shifting landscape of high-performance computing hardware and software. A critical aspect of this adaptation is the development of algorithms that go beyond simply exploiting the brute force of massive parallelism provided by GPUs. While GPUs offer significant acceleration, the impressive performance gain is limited to algorithms using matrix-matrix multiplies with reduced precision. Designing alternative algorithms for QM simulations leveraging this particular strength of GPUs is a must. Furthermore, given the complexity of adapting legacy codes to hybrid architectures, the community must unite in embracing and standardizing modular libraries for quantum chemical calculations, providing reusable, optimized, and versatile implementations of complex algorithms [92]. Some specialized libraries have already demonstrated significant success in enabling efficient implementations on diverse hardware platforms, and their continued development and community-wide adoption are crucial for simplifying code development and ensuring long-term code sustainability. Finally, AI has already proven its broad potential, generating machine-learning force fields with ab initio accuracy, faithfully describing quantum states, and speeding up various tasks within quantum simulations [93]. As a result, the growing integration of AI with quantum simulations offers a promising path for future breakthroughs.
@@
However, a few notes of caution are necessary regarding the widespread adoption of GPUs. CPUs remain essential for workloads that require flexibility, general-purpose computation, or that are memory-bound. For example, CPUs equipped with HBM are competitive with GPUs in some cases, particularly in memory-limited tasks. Although in the near future, we can expect CPUs to integrate AI acceleration features, they will likely remain complementary to GPUs. In addition, while their computational power is undeniable, the energy consumption of GPUs is a growing concern. While a laptop requires less than 100W to function, the peak energy consumption of a modern supercomputer like Frontier is around 22MW. The community must prioritize energy efficiency as a core design principle for future exascale simulations. Time-to-solution should be scaled by power usage when comparing simulations on different architectures. Re-evaluating the blind pursuit of raw performance at the cost of enormous energy consumption (cf. Figs. 1(1(b)) and 4(4(b))) is of paramount importance. The focus should shift towards a more balanced approach where algorithmic efficiency, software optimization, and energy-aware hardware design work in synergy.
@@
Importantly, while the exascale era offers exciting prospects for quantum mechanical simulations, it also demands a cohesive effort from the scientific community to overcome the associated challenges. Collaborative initiatives, such as those driven by global research collaborations and industry partnerships, will be vital in shaping the future of computational chemistry and materials science. By embracing energy-efficient computing, stronger collaboration between academia and industry, and prioritizing the development of standardized software tools (e.g. languages, compilers, debuggers, libraries), we can ensure that exascale computing delivers on its promise as a cornerstone of scientific innovation, enabling us to take on grand challenges and push the boundaries of our understanding of the quantum world.
@@
Acknowledgments The authors acknowledge partial support from the European Centre of Excellence in Exascale Computing TREX — Targeting Real Chemical Accuracy at the Exascale. This project has received funding in part from the European Union’s Horizon 2020 — Research and Innovation Program — under grant agreement no. 952165.
Declarations
Author contribution: R.S. researched data for the article. All authors contributed to the writing of the manuscript.

### v2

@@ Abstract
The exascale era has reinforced the need for a more unified and standardized software stack for GPU programming. Fragmentation creates barriers to code portability, hinders performance optimization, and increases the development effort required for scientific applications. As the community moves forward, it must address these challenges. Initiatives like SYCL, OpenMP, and community-driven efforts like Kokkos and RAJA look promising toward achieving a more unified programming experience. However, vendors, developers, and researchers must collaborate closely to establish common standards and standardize a more cohesive and interoperable software ecosystem for GPU computing.
[H1] QM and HPC
@@
Each QM method relies on distinct computational techniques (such as FFTs, matrix–vector operations or diagonalization of sparse or dense matrices). The techniques determine the scaling behaviour, whereas the specific hardware implementation governs the prefactor associated with this scaling. For instance, matrix multiplication kernels are central to coupled cluster methods that utilize large tensors, enabling efficient use of tensor cores and resulting in a relatively small prefactor despite the high scaling complexity. Conversely, although QMC methods are inherently massively parallel, they often operate on relatively small matrices, leading to a large prefactor; therefore, scaling these approaches on accelerators requires substantial effort [79].
Advances in computational chemistry have increasingly used GPUs to accelerate a range of quantum chemical methods. For Hartree–Fock and DFT calculations, hybrid CPU/GPU implementations have demonstrated notable speedups. Early efforts in this area, such as the development of DFT calculations on multicore and multithreaded hybrid CPU/GPU approaches to Hartree–Fock [130, 131], were pioneering for their time but relied on significantly less powerful GPU hardware than current-generation GPUs. More recent implementations taking advantage of modern GPU architectures have indeed demonstrated even greater performance gains [132, 133]. GPU-accelerated self-consistent field methods have been optimized for large-scale molecular simulations [134, 135, 136]. Coupled-cluster methods, particularly in their density-fitted and non-iterative forms, have also been adapted for GPU architectures to address both single- and multi-reference systems [137, 138, 139, 140, 28, 141]. Additionally, divide-and-conquer approaches have been employed to scale electronic structure calculations for large systems [142, 143, 144]. Real-space grid-based methods offer interesting advantages as they are FFT-free (requiring fewer message passing interface (MPI) communications) and have local operations, and therefore can be massively parallelized.
Beyond traditional methods, GPU-accelerated phaseless auxiliary-field quantum Monte Carlo [145] and multi-GPU configuration interaction [146] techniques have expanded the computational toolkit for tackling strongly correlated systems. Finally, specific applications, such as accelerating plane-wave-based calculations in VASP [147] and implementing GPU-based GW calculations [148], further highlight the versatility of GPU acceleration in computational chemistry.
The efficient evaluation of electron repulsion integrals on GPUs has become a central focus in molecular quantum chemistry. Developments in this area offer a great deal of potential to accelerate electronic structure calculations. Early foundational efforts [149, 150] introduced GPU-based approaches to integral evaluation, which have since been refined through a variety of algorithms and optimization strategies, including the use of mixed-precision techniques [151, 152, 153]. Subsequent developments have addressed challenges such as higher angular momentum integrals [154, 155, 156], algorithmic innovations like the BRUSH algorithm [157], and automated code generation for GPU-based integral evaluation [158]. Hybrid CPU/GPU approaches [159] and advancements in GPU compiler technologies [160] have further expanded the computational possibilities. The use of multinode, multi-GPU systems for large-scale calculations has been recently demonstrated [161], as well as high-performance integral evaluation implementations [162].
The adoption of mixed single and double precision has been explored at various stages in computational methods. Initially, this approach was investigated at the level of linear system solvers [163]. Subsequently, it gained traction in quantum chemistry methods, including MP2 [164, 165, 166], QMC simulations [167], coupled-cluster methods [168] and quantum molecular dynamics [169]. The integration of single precision was first introduced in early GPU-accelerated quantum chemistry codes, notably through optimized matrix multiplication libraries [165] and then in the acceleration of two-electron integral computations [151, 153, 152, 131]. These implementations provided substantial speedups by leveraging the performance advantages of single-precision arithmetic on GPUs. The use of mixed precision for the solution of the non-linear Kohn–Sham eigenvalue problem in the DFT-FE code resulted in the execution time being reduced by approximately a factor of two compared to double precision for a massively parallel calculation [65]. Furthermore, the use of single precision in initial iterations of eigensolvers speeds up calculations on a massive scale without compromising final accuracy [170].
More recently, research has extended to the use of half-precision (16-bit floating point) as a means of further optimizing computational performance while maintaining an acceptable level of accuracy. Its potential benefits and challenges have been explored in real applications [171, 172], including in the context of quantum chemistry [173], where accurate approximations for matrix products using fast and error-free splitting of floating-point numbers [174] were explored. Mixed precision strategies in the WEST code achieved significant speedups for large-scale, full-frequency GW calculations [175]. Furthermore, calculations of exchange interactions in periodic systems using single-precision arithmetic, particularly within hybrid DFT, have demonstrated substantial reductions in computation time and memory requirements with negligible loss of accuracy in band energies, forces, and x-ray absorption spectra [176].
@@
Efforts to adapt to the heterogeneous architectures of HPC systems highlight the critical importance of balancing performance with portability. Some codes, such as ABINIT and Quantum Espresso, have adopted hardware-agnostic offload programming models (OpenMP or OpenACC), whereas others, such as CP2K, have opted for non-portable, kernel-based models (CUDA and HIP). BerkeleyGW provides both CUDA and OpenACC alternatives, with the latter achieving performance within ten percent of the former [148]. Additionally, different communities are actively developing specialized performance libraries to execute computationally intensive tasks on accelerators. Examples include the Distributed Block Compressed Sparse Row Matrix (DBCSR) library [177] for Nvidia and AMD GPUs and the LibintX [178, 179] library, which provides efficient implementations of Gaussian integral evaluation, central to integral-direct implementations. The QM community library projects such as ELSI [180], and especially ELPA, have proved very efficient and have been adopted by many codes[181]. The QMCkl library, developed within the TREX CoE, offers high-performance computation kernels for QMC simulations using OpenMP and OpenACC [26]. The wave-function community [68, 29, 48, 182, 183, 184] has been extensively developing tensor libraries that automatically parallelize and accelerate operations on multidimensional arrays, taking advantage of modern HPC architectures [185, 186, 187].
Although many studies have reported improved timings for QM calculations on heterogeneous architectures, the energy-to-solution costs remain a relatively unexplored aspect. Also, in this regard a direct comparison of timings with CPU-only systems can be misleading because these architectures differ significantly in power consumption, memory bandwidth, and hardware complexity, all of which should be accounted for. For example, an energy-efficient distributed, block-cyclic setup of the Hamiltonian and overlap matrices for linearized augmented plane wave calculations has been reported [188]. Another large-scale benchmark study on the Summit supercomputer — the nodes of which house two IBM POWER9 CPU sockets and six NVIDIA V100 GPUs connected via high-bandwidth NVLink — further underscores the architectural disparities [90]. When normalized under equivalent power usage, the GPU-accelerated system provided up to sevenfold speedups compared to CPU-only (that is, using only the CPUs from these heterogeneous nodes) system.
Supercomputer | Year | Configuration | Power | HPL | HPCG |
(MW) | (Pflops) | (Pflops) | |||
Fugaku | 2020 | CPU-only | |||
Frontier | 2021 | CPU/GPU | |||
Aurora | 2024 | CPU/GPU |
However, it is crucial to interpret such results with caution. This type of comparison, while isolating the contribution of the GPU, does not represent a realistic operational scenario. Furthermore, measuring the speedup of a GPU should not be done by comparison with the CPUs hosting those GPUs, but by a direct comparison to a separate, dedicated CPU-only system, which typically employs better-performing, higher-end CPUs. Therefore, these results illustrate that although heterogeneous systems can deliver remarkable performance gains, their inherently different hardware profiles preclude an entirely fair one-to-one comparison with CPU-only configurations without simultaneously accounting for power requirements, memory constraints and scaling behaviour (see Table 3).
[H1] Outlook
@@
The future of QM simulations at the exascale level hinges on effectively navigating the shifting landscape of HPC hardware and software. A critical aspect of this adaptation is the development of algorithms that go beyond simply exploiting the brute force of massive parallelism provided by GPUs. Although GPUs offer substantial acceleration, the impressive performance gain is limited to algorithms that use matrix–matrix multiplication with reduced precision. Designing alternative algorithms for QM simulations that make use of this particular strength of GPUs is a must. Furthermore, given the complexity of adapting legacy codes to hybrid architectures, the community must unite in embracing and standardizing modular libraries for quantum chemical calculations, providing reusable, optimized and versatile implementations of complex algorithms [27]. Some specialized libraries have already succeded significant success in enabling efficient implementations on diverse hardware platforms, and their continued development and community-wide adoption are crucial for simplifying code development and ensuring long-term code sustainability. Finally, AI has already proven its broad potential, generating machine-learning force fields with ab initio accuracy, faithfully describing quantum states and speeding up various tasks within quantum simulations [189]. As a result, the growing integration of AI with quantum simulations offers a promising path for future breakthroughs.
@@
However, a few notes of caution are necessary regarding the widespread adoption of GPUs. CPUs remain essential for workloads that require flexibility, general-purpose computation, or that are memory-bound. For example, CPUs equipped with HBM are competitive with GPUs in some cases, particularly in memory-limited tasks. Although in the near future, we can expect CPUs to integrate AI acceleration features, they will likely remain complementary to GPUs. In addition, although their computational power is undeniable, the energy consumption of GPUs is a growing concern. Whereas a laptop requires less than 100W to function, the peak energy consumption of a modern supercomputer like El Capitan is around 29MW. This stark contrast, while not intended as a direct performance comparison, highlights the immense difference in scale and the critical need for energy-efficient exascale computing. Even though the total energy consumption of exascale systems will inevitably be higher than previous generations due to their increased scale and capabilities, the community must prioritize energy efficiency as one of the central design principles for future exascale simulations. Time-to-solution should be accompanied by energy-to-solution when comparing simulations on different architectures. Re-evaluating the blind pursuit of raw performance at the cost of enormous energy consumption is of paramount importance. The focus should shift towards a more balanced approach in which algorithmic efficiency, software optimization and energy-aware hardware design work in synergy.
@@
Importantly, although the exascale era offers exciting prospects for QM simulations, it also demands a cohesive effort from the scientific community to overcome the associated challenges. Collaborative initiatives, such as those driven by global research collaborations and industry partnerships, will be vital in shaping the future of computational chemistry and materials science. By embracing energy-efficient computing, stronger collaboration between academia and industry, and prioritizing the development of standardized software tools (such as languages, compilers, debuggers and libraries), we can ensure that exascale computing delivers on its promise as a cornerstone of scientific innovation, enabling us to take on grand challenges and push the boundaries of our understanding of the quantum world.
Acknowledgments
@@
The authors acknowledge partial support from the European Centre of Excellence in Exascale Computing TREX — Targeting Real Chemical Accuracy at the Exascale. This project has received funding in part from the European Union’s Horizon 2020 — Research and Innovation Program — under grant agreement no. 952165.
Author contributions
All authors contributed to all aspects of this work.
Competing interests
The authors declare no competing interests.